given that loga2= 0.693, and loga11= 2.398, find loga 1/22

given that log x = 3 and log y = 6, find log xy2

Answer by oscargut(686)

given that loga2= 0.693, and loga11= 2.398, find loga 1/22
loga(1/22)=loga(1)-loga(22)=0-loga(2*11)=-(loga(2)+loga(11))=-(0.693+2.398)=-3.091

given that log x = 3 and log y = 6, find log xy2
log(xy^2)=log(x)+log(y^2)=log(x)+2log(y)=3+2(6)=15